OpenStudy (gtxmuqsit):

A 40 Kg block is resting at a height of 5 m off the ground. If the block is released and falls to the ground, what is its total energy at a height of 2m? (g=10m/s2). A) 0J B) 400J C) 2J D) 6J E) cannot be determined from given info. @AllTehMaffs

OpenStudy (anonymous):

right? @GTXMUQSIT

OpenStudy (anonymous):

\[E = K + U\] \[k = \frac{ 1 }{ 2 }mv^{2}\] \[U = mgh\] Where m = mass, v = velocity, h = height.

OpenStudy (anonymous):

The total energy at any given point is given by that relation (E = K + U). At the height of 5 meters, where the object is resting. The total kinetic energy is 0, because v=0. So E = K(5m) + U(5m) = U(5m)

OpenStudy (anonymous):

At the height of 2 meters. E = K(2m) + U(2m) = U(5m) You can calculate velocity using freefall equations.

OpenStudy (anonymous):

but the total energy at that point has to be exactly the energy as when it strted falling

OpenStudy (anonymous):

Yeah, but |mgh| = |(40(10)(5)| = 2000. Which is way to big. I think it's asking for K(2m) or U(2m).

OpenStudy (anonymous):

God I need to learn how to multiply

OpenStudy (anonymous):

I think something is missing. It doesn't add up.

OpenStudy (anonymous):

(Neither K or U)

OpenStudy (anonymous):

blergh

OpenStudy (gtxmuqsit):

@Lessis and @AllTehMaffs sir I couldn't find answer I apply the following formula P.E=mgh so P.E=40x10x2= 800 J which is not in options so I think option E is correct

OpenStudy (gtxmuqsit):

@sarah786 do you know the anwser?

OpenStudy (gtxmuqsit):

@CGGURUMANJUNATH

OpenStudy (gtxmuqsit):

would you help

OpenStudy (anonymous):

I think the answer isn't given - none of them make sense. It asks for "total energy," which can't be any of the options. \[ E_{initial}=E_{tot}=40kg*10m/s^2*5m = 2000J\] Potential energy at 2m does not give the correct answer (like you saw) and \[K = E_{tot} - U_{2m} = 2000J-(40*10*2)=1200J\] so that's not the right answer either. It has to be E - there *is* enough information given, they just don't give the correct answer for you to choose ^_^

OpenStudy (gtxmuqsit):

@AllTehMaffs thank you sir :D

OpenStudy (anonymous):

maybe those are KJ in which case C is the answer :P